OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (5/3,-7) and has a slope of 3/4

OpenStudy (anonymous):

do -7=3/4*5/3+b solve for b

OpenStudy (anonymous):

b=-8.25

OpenStudy (anonymous):

so y=3/4x-8.2

OpenStudy (anonymous):

in fractions?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

y=3/4x-41/5 is the answer on fractions